write an equation of the line containing the given point and parallel to the given line. Express your answer in the form y=mx+b.
(2,6); x+6y=5, the equation of the line is y=?

Answer by ewatrrr(24785)


Hi
Line passing thru Pt(2,6) and parallel to x+6y=5


x+6y=5 OR y = (-1/6)x + 5/6 |slope of these parallel lines is m = -1/6
New Line
y = (-1/6)x + b |using ordered pair Pt(2,6) to solve for b
6 = -1/3 + b
18/3 + 1/3 = b
19/3 = b
y = (-1/6)x + 19/3 |y-intercept is Pt(0,19/3)